OpenStudy (anonymous):

sec(pi/6) exact value

OpenStudy (anonymous):

sec(pi/6) = 1/cos(pi/6) = 1/(sqrt(3)/2) = \[\frac{2\sqrt{3}}{3}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem :)